# METok: Multi-Stage Event-based Token Compression for Efficient Long Video Understanding

### v1

@@ Appendix D More Details of Experiments
Notably, as the number of frames increases, METok not only maintains performance but also brings consistent gains, especially on the MLVU and VideoMME benchmarks. For instance, at 64 frames, METok improves the MLVU score from 50.4 to 53.6, and VideoMME from 49.5 to 50.7. These results indicate that METok effectively compresses less informative tokens while preserving task-relevant visual semantics, allowing the model to benefit from richer temporal context.
@@
$\alpha$ $\beta$ FLOPs (T) Egoschema VideoMME Base Model (32 frames) 71.4 43.5 58.2 0.3 0.2 14.6 42.8 57.3 0.3 0.4 17.9 43.3 57.6 0.5 0.2 17.2 43.7 58.2 0.5 0.4 21.8 44.4 58.4 0.7 0.2 34.5 44.3 58.2 0.7 0.4 43.3 44.4 58.7
@@ Appendix E Additional Qualitative Examples
We provide additional video question answering examples from video understanding benchmarks like MLVU, as shown in Figure 9 and Figure 10. These examples further demonstrate the effectiveness of the proposed METok framework in capturing key visual-semantic cues from long videos. METok consistently produces precise and contextually grounded answers, successfully identifying fine-grained actions, object interactions, and temporal dependencies. By focusing on the relevant visual segments and filtering out redundant frames, METok generates accurate and coherent responses.

### v2

@@ Appendix D More Details of Experiments
Notably, as the number of frames increases, METok not only maintains performance but also brings consistent gains, especially on the MLVU and VideoMME benchmarks. For instance, at 64 frames, METok improves the MLVU score from 50.4 to 53.6, and VideoMME from 49.5 to 50.7. These results indicate that METok consistently achieves strong accuracy across different frame sampling densities and effectively compresses less informative tokens while preserving task-relevant visual semantics, allowing the model to benefit from richer temporal context.
@@
Below, we further compare different token reduction methods with VILA-1.5-13B under the 32-frame inputs. As shown in Table 7, METok achieves the lowest FLOPs while outperforming others in MVBench, MLVU, and VideoMME benchmarks. Notably, VILA-1.5-13B has a native maximum context length of 4096, which limits it to processing only 16 frames without any token reduction. Within this controlled setting, METok delivers better trade-off between accuracy and computation, indicating that its multi-stage, event-aware compression preserves task-relevant cues while minimizing redundant computation.
@@
Method FLOPs (TB) MVBench MLVU VideoMME w/FastV 78.0 50.4 52.0 48.9 w/VisionZip 56.7 48.9 50.3 47.9 w/DivPrune 37.2 50.9 52.5 49.4 w/METok 36.5 50.9 52.8 49.9
Table 7: Comparison of token-reduction methods with VILA-1.5-13B under 32-frame inputs.
Method FLOPs (TB) AR NQA TR PQA AO AC ER Overall
VILA-1.5-13B 84.1 52.5 53.0 78.3 57.1 35.9 23.6 52.0 50.4
w/FastV 49.8 52.0 52.4 78.7 56.4 35.0 22.0 50.0 49.5
w/VisionZip 29.1 46.5 51.3 72.2 50.7 36.9 22.0 50.0 47.1
w/DivPrune 20.3 51.4 54.6 78.7 57.2 35.8 22.2 51.9 50.3
w/METok (Ours) 19.6 50.6 55.3 78.8 57.8 36.9 22.5 51.3 50.5
Table 8: Results on MLVU subsets with VILA-1.5-13B.
@@
$\alpha$ $\beta$ FLOPs (TB) Egoschema VideoMME Base Model (32 frames) 71.4 43.5 58.2 0.3 0.2 14.6 42.8 57.3 0.3 0.4 17.9 43.3 57.6 0.5 0.2 17.2 43.7 58.2 0.5 0.4 21.8 44.4 58.4 0.7 0.2 34.5 44.3 58.2 0.7 0.4 43.3 44.4 58.7
@@
D.3 Performance across MLVU subsets
To provide a more comprehensive evaluation, we conducted a detailed analysis across all seven MLVU subcategories. As shown in Table 8, METok achieves the highest or competitive accuracy in most task types and consistently outperforms all baselines in overall accuracy and FLOPs reduction (from 84.1 to 19.6 TB). It particularly excels in long-context reasoning tasks such as NQA and PQA, where its token compression effectively preserves key visual-text semantics over time. These results not only demonstrate the task-level robustness and generality of METok, but also highlight its strength in long-form video question answering.
@@ Appendix E Additional Qualitative Examples
We provide additional video question answering examples from video understanding benchmarks like MLVU, as shown in Figure 9 and Figure 10. In the background recognition example, METok with LongVA-7B correctly selects “Windmills” as the scene behind the engineer working with drawings, indicating robust scene-level understanding under distractor options. In the object retrieval case from an animated domain, METok accurately identifies that the cartoon lobster lifts paper money, showing resilience to style shifts and the ability to ground fine object semantics. For an attribute recognition query, METok answers white for the flower color, demonstrating precise grounding of low-level visual attributes. Together, these cases span scene context, object semantics, and fine-grained attributes, and they align with our design goal of preserving key visual-text cues while suppressing redundancy in long videos.
These examples further demonstrate the effectiveness of the proposed METok framework in capturing key visual-semantic cues from long videos. METok consistently produces precise and contextually grounded answers, successfully identifying fine-grained actions, object interactions, and temporal dependencies. By focusing on the relevant visual segments and filtering out redundant frames, METok generates accurate and coherent responses.